OpenStudy (anonymous):

Solve.3 over x = 1 over x minus 2 A. x = 3 B. x = 4 C. x = 6 i get 2 can someone tell me where im going wrong

OpenStudy (anonymous):

1/x - 2?

OpenStudy (anonymous):

3/x=1/x-2

OpenStudy (anonymous):

you cross multiply

OpenStudy (anonymous):

i did and then divided by 3 and got x=2

OpenStudy (anonymous):

1(x)=3(x-2)

OpenStudy (anonymous):

but its not one of the answers

OpenStudy (anonymous):

oh okay i see where i went wrong

OpenStudy (anonymous):

1(3)=3(3-2) 3=9-6 3=3 3-3=0 0=0 Thats what i got for a

OpenStudy (anonymous):

thats what i got too